News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Monday 17 April 2000 11.00am EDT

During the last week, Chandra completed the observing schedule as planned.

A real-time procedure was run at 5.30pm EDT on 3/15 with ACIS to perform the second Squeegee-mode test designed to improve the energy resolution of the degraded chips. The test was conducted with the S0 chip and included runs with an optimized 16 rows, 8 rows and a control run. The test was completed successfully and initial results indicate an improvement due to the optimization. Several more tests are planned over the next several weeks to determine the optimal parameters for the new mode.

Other real-time procedures run during the week included a routine EIA Sequencer self-check and a routine diagnostic dump of the CPE warm start counter.

The planned observations for next week are as follows.

NGC 4395               ACIS-S          Apr 17
NGC 4258               ACIS-S
Jet of 3C 390.3        ACIS-S
Radiation Belts
XTE J1118+48           ACIS-S/LETG     Apr 18
Mrk 273X               ACIS-S
RX J1720.1+2638        ACIS-I          Apr 19
H1413+117              ACIS-S
NGC 4649               ACIS-S
NGC 4278               ACIS-S          Apr 20
M87                    ACIS-S
NGC 4374               ACIS-S
Radiation Belts
V382Velorum1999        ACIS-I          Apr 21
EMSS 1054.5-0321       ACIS-S
NGC 2110               ACIS-S          Apr 22
Abell 3125             ACIS-I
NGC 526A               ACIS-S          Apr 23
Radiation Belts
GRS 1915+105           ACIS-S/HETG


The spacecraft subsystems continued to operate nominally.

Return to Status Reports